How many electrons must pass through a circuit to account for 106 coulombs of charge?
1
Expert's answer
2012-09-07T10:22:27-0400
Total charge Q = 106 C.
The charge of single electron is 1.6 * 10^-19 C.
N
= Q / e = 106 / (1.6 * 10^-19) = 6.625 * 10^20.